# Affine Chabauty I

### v1

@@ 1. Introduction
an inequality involving the genus $g$, the Mordell–Weil rank $r$, the cardinality of $S$, and a contribution from the boundary $D$ of $Y$. Our method is an $S$-integral analogue of the method of Chabauty–Coleman [Cha41, Col85]. Here, we embed $Y$ into its generalised Jacobian $J_{Y}$, choose an auxiliary prime $p\not\in S$ (possibly of bad reduction), and cut out $\mathcal{Y}(\mathbb{Z}_{S})$ inside $\mathcal{Y}(\mathbb{Z}_{p})$ by $p$-adic integrals of logarithmic differentials, i.e. meromorphic differentials on a compactification of $Y$ with at worst simple poles at $D$. The existence of suitable log differentials is guaranteed by assumption (1.1), which ensures that the image of $\mathcal{Y}(\mathbb{Z}_{S})$ inside $J_{Y}(\mathbb{Q})$ is contained in a finite union of translates of subgroups of controllable rank. We then bound the size of $\mathcal{Y}(\mathbb{Z}_{S})$ by estimating the number of zeros of $p$-adic integrals of log differentials. In a follow-up article [LL], we present an algorithm to determine the log differentials and the zeros of their integrals in practice, thus computing a finite subset of $\mathcal{Y}(\mathbb{Z}_{p})$ containing $\mathcal{Y}(\mathbb{Z}_{S})$.
@@ 1.1. Main results
In this introduction, we work with curves over $\mathbb{Q}$, but in the body of the paper we formulate our results over number fields. Let $Y/\mathbb{Q}$ be given as $X\smallsetminus D$ where $X/\mathbb{Q}$ is a smooth projective curve and $D\neq\emptyset$ is a finite set of closed points called cusps. Let $S$ be a finite set of primes. Let $\mathcal{X}$ be a regular model of $X$ [Liu06, Definition 10.1.1] over the ring $\mathbb{Z}_{S}$ of $S$-integers. Let $\operatorname{\mathcal{D}}$ be the closure of $D$ in $\mathcal{X}$ and set $\mathcal{Y}\coloneqq\mathcal{X}\smallsetminus\operatorname{\mathcal{D}}$. Fix a base point $P_{0}\in Y(\mathbb{Q})$ and a prime $p\not\in S$. We use the following notation, which will be kept throughout this paper:
@@
Let $\Sigma$ be an $S$-integral reduction type and let $C(\Sigma)\subseteq S$ denote the set of those $\ell\in S$ where $\Sigma_{\ell}$ is cuspidal, compare ˜3.10. Let $p\not\in S$ be a prime. Assume
@@
Using approach (2), we previously showed in joint work with Steffen Müller that hypothesis (1.1) implies finiteness of the depth-1 Chabauty–Kim locus [LLM23, Theorem A(1)+Remark 1.4], and obtained bounds under a more restrictive hypothesis [LLM23, Remark 1.2]. The present manuscript improves upon this in several aspects: the bound in ˜C is smaller, we only need hypothesis (1.1) to prove both finiteness and a bound, we allow $p$ to be a prime of bad reduction for $\mathcal{X}$, our method works over general number fields, and we determine the type of Coleman functions on $\mathcal{Y}(\mathbb{Z}_{p})$ vanishing on $S$-integral points. Crucial for these improvements is the geometry and arithmetic of the generalised Jacobian. Moreover, the functions $\rho$ in ˜A can be algorithmically determined in practice, and by computing their zeros one finds a finite subset of $\mathcal{Y}(\mathbb{Z}_{p})$ containing $\mathcal{Y}(\mathbb{Z}_{S})$. We will present the computational aspects of our method in a follow-up article [LL].
Our method here can also be seen as a refined version of (3) in the sense of [BD20]. Namely, the idea of (3) is to embed $\mathcal{Y}$ into a suitable model $\mathcal{J}/\mathbb{Z}_{S}$ of the generalised Jacobian $J_{Y}$. Finiteness of $\mathcal{Y}(\mathbb{Z}_{S})$ is ensured by
@@ 1.3. Structure of the paper
In Section˜2, we define the $D$-intersection map $\sigma_{\mathfrak{m}}$ in a local setting, namely for affine curves over discretely valued fields. We introduce $D$-transversal models, integral and rational reduction types $\Sigma_{\mathfrak{m}}$ and show that the composition $\sigma_{\mathfrak{m}}\circ\operatorname{AJ}_{P_{0}}$ restricted to points of a given reduction type lands in the subset $\mathfrak{S}_{\mathfrak{m}}(P_{0},\Sigma_{\mathfrak{m}})$. We also compare the $D$-intersection map with the local Kummer map. In Section˜3, we define the global $D$-intersection map $\sigma$ for an affine curve over a number field, and we determine the rank of $\ker(\sigma)$. We define $S$-integral reduction types $\Sigma$ and the global Selmer set $\operatorname{Sel}(P_{0},\Sigma)$. We then show that the Abel–Jacobi images of $S$-integral points of reduction type $\Sigma$ are contained in $\operatorname{Sel}(P_{0},\Sigma)$ and bound the rank of $\operatorname{Sel}(P_{0},\Sigma)$. In Section˜4, we summarise the theory of $p$-adic abelian integrals of log differentials. In Section˜5 we prove ˜A and describe an $S$-integral version of Siksek’s Restriction of Scalars Chabauty [Sik13]. We then proceed to prove ˜C in Section˜6, and end this paper in Section˜7 with two applications of our methods to even degree hyperelliptic curves: an improved bound on the number of integral points of such curves, and the complete determination of the $\mathbb{Z}[\zeta_{3}]$-points of a curve of genus 2 and Mordell–Weil rank 2.
@@ Acknowledgements
We thank Jennifer Balakrishnan, Alex Betts, Stevan Gajović, Sachi Hashimoto, Enis Kaya, Steffen Müller, Bjorn Poonen, Pim Spelier, and Elie Studnia for helpful discussions on this paper. The first author is supported by a Walter Benjamin Scholarship from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), project number LE 5634/1, and also acknowledges support from the DFG through TRR 326 Geometry and Arithmetic of Uniformized Structures, project number 444845124. The second author is supported by a Minerva Fellowship of the Minerva Stiftung Gesellschaft für die Forschung mbH and also acknowledges support through a guest postdoc fellowship at the Max Planck Institute for Mathematics in Bonn and through an NWO Grant, project number VI.Vidi.192.106.
@@ 2.2. Intersection numbers
For the remainder of this section, let $(R,\mathfrak{m},k)$ be a discrete valuation ring with field of fractions $K$, let $X$ be a smooth proper curve over $K$, let $D\neq\emptyset$ be a finite set of closed points of $X$ and set $Y=X\smallsetminus D$. In order to construct the $D$-intersection map $\sigma_{\mathfrak{m}}\colon J_{Y}(K)\to V_{D,\mathfrak{m}}$, we choose a regular model $\mathcal{X}$ of $X$ over $R$, i.e. an integral, regular, projective and flat $R$-scheme with an isomorphism $\mathcal{X}_{K}\cong X$ [Liu06, Definition 10.1.1]. We will show later that the map $\sigma_{\mathfrak{m}}$ does not depend on the chosen model. Recall from [Lan88, §III.2] the definition of intersection numbers on the arithmetic surface $\mathcal{X}$. If $E,F$ are effective divisors on $\mathcal{X}$ without common component and $x$ is a closed point of $\mathcal{X}$, the intersection number of $E$ and $F$ at $x$ is defined as
@@
$i_{\mathfrak{m}}(E,F)\coloneqq E\cdot F\coloneqq\deg_{k}(E.F)=\sum_{x}i_{x}(E,F)[k(x):k].$
By replacing a divisor with a linearly equivalent one, one can also define the self-intersection number $i_{\mathfrak{m}}(E,E)$ of a vertical prime divisor $E$. Hence, the symbol $i_{\mathfrak{m}}$ extends to all pairs of divisors on $\mathcal{X}$ with no horizontal common component [Lan88, §III.3].
@@ 2.6. $D$-transversal models
As $R$ is excellent, by [Liu06, Theorem 9.2.26] successive blow-ups of $\mathcal{X}$ in closed points [^1] As we only need $\operatorname{\mathcal{D}}$ to be regular at all $x\in\mathcal{X}_{\mathfrak{m}}^{\mathrm{sm}}(k)$, blow-ups in those points $x\in(\mathcal{X}_{\mathfrak{m}}^{\mathrm{sm}}\cap\operatorname{\mathcal{D}})(k)$ with $i_{x}(\mathcal{X}_{\mathfrak{m}},\operatorname{\mathcal{D}})\geq 2$ should suffice, but we did not check whether the analogue of [Liu06, Lemma 9.2.32] is true. yield a regular model $\mathcal{X}^{\prime}$ of $X$ in which the closure $\operatorname{\mathcal{D}}^{\prime}$ of $D$ is normal. Now we blow up again in every $x\in(\mathcal{X}_{\mathfrak{m}}^{\prime,\mathrm{sm}}\cap\operatorname{\mathcal{D}}^{\prime})(k)$ such that $i_{x}(\mathcal{X}^{\prime}_{\mathfrak{m}},\operatorname{\mathcal{D}}^{\prime})\geq 2$ to obtain the model $\mathcal{X}^{\prime\prime}$. We claim that $\mathcal{X}^{\prime\prime}$ is a $D$-transversal regular model of $X$. Indeed, away from those $x$ the model $\mathcal{X}^{\prime}$ is already $D$-transversal. For $x\in(\mathcal{X}_{\mathfrak{m}}^{\prime,\mathrm{sm}}\cap\operatorname{\mathcal{D}}^{\prime})(k)$, let $Q$ be the unique cusp such that $x$ lies in the closure $\mathcal{Q}^{\prime}$ of $Q$. If $2\leq i_{x}(\mathcal{X}^{\prime}_{\mathfrak{m}},\operatorname{\mathcal{D}}^{\prime})=i_{x}(\mathcal{X}^{\prime}_{\mathfrak{m}},\mathcal{Q}^{\prime})$, let $E$ denote the exceptional divisor on $\mathcal{X}^{\prime\prime}$ of the blow-up in $x$. Then $\mathcal{Q}^{\prime\prime}$ intersects $\mathcal{X}^{\prime\prime}_{\mathfrak{m}}$ in the intersection point of $E$ with the strict transform of $\mathcal{X}_{\mathfrak{m}}^{\prime,\mathrm{sm}}$, which is not a point in $\mathcal{X}_{\mathfrak{m}}^{\prime\prime,\mathrm{sm}}$, so there is no $D$-transversality left to check. ∎
@@ 3. Global theory: Selmer subsets of the generalised Jacobian
Let $K$ be a number field, let $X$ be a smooth proper curve over $K$, let $D\neq\emptyset$ be a finite set of closed points of $X$ and set $Y=X\smallsetminus D$. The ring of integers of $K$ is denoted by $\mathcal{O}_{K}$. We start by defining a $\mathbb{Q}$-vector space $V_{D}$ and a homomorphism $\sigma\colon J_{Y}(K)\to V_{D}$. In Section˜3.2 we define $S$-integral reduction types with respect to a regular model over $\mathcal{O}_{K}$ and use the map $\sigma$ to constrain the image of points of a fixed reduction type under an Abel–Jacobi map $Y\to J_{Y}$.
@@ 3.3. Selmer sets
Denote by $C(\Sigma)$ the set of primes $\mathfrak{q}\in S$ for which $\Sigma_{\mathfrak{q}}$ is cuspidal, i.e. an element of $\mathcal{C}_{\mathfrak{q}}^{1}=(\mathcal{X}_{\mathfrak{q}}^{\mathrm{sm}}\cap\operatorname{\mathcal{D}})(\operatorname{\mathbb{F}}_{\mathfrak{q}})$.
@@
Assume that $D$ does not consist of a single $K$-rational point. Then $\mathfrak{S}(P_{0},\Sigma)$ is a translate of a subgroup of $V_{D}$ which is free of rank $C(\Sigma)$.
@@
We compare the Selmer set $\operatorname{Sel}(P_{0},\Sigma)\subseteq J_{Y}(K)$ with the cohomological Selmer set $\operatorname{Sel}_{\Sigma,U_{1}}\subseteq\mathrm{H}^{1}(G_{K},V_{p}J_{Y})$ of [Bet23, §6.2], where $U_{1}=V_{p}J_{Y}$ is the abelianisation of the $\mathbb{Q}_{p}$-prounipotent fundamental group of $Y$. Assume that all primes $\mathfrak{q}|p$ are primes of good reduction for $(\mathcal{X},\operatorname{\mathcal{D}})$, that $P_{0}$ is integral at those primes, and that they are not contained in $S$. The set $\operatorname{Sel}_{\Sigma,U_{1}}$ is defined as the set of those global cohomology classes whose restriction at primes $\mathfrak{q}|p$ is crystalline, and whose restriction lies in $\mathfrak{S}^{\mathrm{coh}}_{\mathfrak{q}}(P_{0},\Sigma_{\mathfrak{q}})\subseteq\mathrm{H}^{1}(G_{K_{\mathfrak{q}}},V_{p}J_{Y})$ for all primes $\mathfrak{q}\nmid p$; see §2.8 for the definition of $\mathfrak{S}_{\mathfrak{q}}^{\mathrm{coh}}(P_{0},\Sigma_{\mathfrak{q}})$. ([Bet23, §6] also assumes that the base point $P_{0}$ is $S$-integral and that $K=\mathbb{Q}$ but the definition makes sense more generally.) Our assumptions ensure that there is only one reduction type $\Sigma_{\mathfrak{q}}$ at primes $\mathfrak{q}\mid p$ and $\mathfrak{S}_{\mathfrak{q}}(P_{0},\Sigma_{\mathfrak{q}})=\{0\}$. Let $\kappa\colon J_{Y}(K)\to\mathrm{H}^{1}(G_{K},V_{p}J_{Y})$ denote the global Kummer map, defined in the same way as in (2.15) and compatible with the local Kummer maps. Now ˜2.33 and ˜2.36 show

### v2

@@ 1. Introduction
an inequality involving the genus $g$, the Mordell–Weil rank $r$, the cardinality of $S$, the number of closed points and the number of conjugate pairs of complex points in the boundary $D$ of $Y$. Our method is an $S$-integral analogue of the method of Chabauty–Coleman [Cha41, Col85]. Here, we embed $Y$ into its generalised Jacobian $J_{Y}$, choose an auxiliary prime $p\not\in S$ (possibly of bad reduction), and cut out $\mathcal{Y}(\mathbb{Z}_{S})$ inside $\mathcal{Y}(\mathbb{Z}_{p})$ by $p$-adic integrals of logarithmic differentials, i.e. meromorphic differentials on a compactification of $Y$ with at worst simple poles at $D$. The existence of suitable log differentials is guaranteed by assumption (1.1), which ensures that the image of $\mathcal{Y}(\mathbb{Z}_{S})$ inside $J_{Y}(\mathbb{Q})$ is contained in a finite union of translates of subgroups of controllable rank. We then bound the size of $\mathcal{Y}(\mathbb{Z}_{S})$ by estimating the number of zeros of $p$-adic integrals of log differentials. In a follow-up article [LL], we present an algorithm to determine the log differentials and the zeros of their integrals in practice, thus computing a finite subset of $\mathcal{Y}(\mathbb{Z}_{p})$ containing $\mathcal{Y}(\mathbb{Z}_{S})$.
@@ 1.1. Main results
In this introduction, we work with curves over $\mathbb{Q}$, but in the body of the paper we formulate our results over number fields. Let $Y/\mathbb{Q}$ be given as $X\smallsetminus D$ where $X/\mathbb{Q}$ is a smooth projective curve and $D\neq\emptyset$ is a finite set of closed points called cusps. Let $S$ be a finite set of primes. Let $\mathcal{X}$ be a regular model of $X$ over the ring $\mathbb{Z}_{S}$ of $S$-integers, i.e., a regular, flat, projective $\mathbb{Z}_{S}$-scheme with an isomorphism $\mathcal{X}_{\mathbb{Q}}\cong X$ [Liu06, Definition 10.1.1]. Let $\operatorname{\mathcal{D}}$ be the closure of $D$ in $\mathcal{X}$ and set $\mathcal{Y}\coloneqq\mathcal{X}\smallsetminus\operatorname{\mathcal{D}}$. We use the following notation, which will be kept throughout this paper:
@@
Let $\Sigma$ be an $S$-integral reduction type and let $C(\Sigma)\subseteq S$ denote the set of those $\ell\in S$ where $\Sigma_{\ell}$ is cuspidal, see ˜3.10. Let $p\not\in S$ be a prime. Assume
@@
Using approach (2), we previously showed in joint work with Steffen Müller that hypothesis (1.1) implies finiteness of the depth-1 Chabauty–Kim locus [LLM23, Theorem A(1)+Remark 1.4], and obtained bounds under a more restrictive hypothesis [LLM23, Remark 1.2]. The present manuscript improves upon this in several aspects: the bound in ˜C is smaller, we only need hypothesis (1.1) to prove both finiteness and a bound, we allow $p$ to be a prime of bad reduction for $\mathcal{X}$, our method works over general number fields, and we determine the type of Coleman functions on $\mathcal{Y}(\mathbb{Z}_{p})$ vanishing on $S$-integral points. Crucial for these improvements is the geometry and arithmetic of the generalised Jacobian. Moreover, the functions $\rho$ in ˜A can be algorithmically determined in practice, and by computing their zeros one finds a finite subset of $\mathcal{Y}(\mathbb{Z}_{p})$ containing $\mathcal{Y}(\mathbb{Z}_{S})$. In the case of an even-degree hyperelliptic curve over $\mathbb{Q}$ and $S=\emptyset$, such a method was recently developed in [GM25] using $p$-adic Coleman–Gross heights. This “Linear Quadratic Chabauty” method arises as a special case of our method, see Section 7.1.
Our method can also be seen as a refined version of (3) in the sense of [BD20]. Namely, the idea of (3) is to embed $\mathcal{Y}$ into a suitable model $\mathcal{J}/\mathbb{Z}_{S}$ of the generalised Jacobian $J_{Y}$. Finiteness of $\mathcal{Y}(\mathbb{Z}_{S})$ is ensured by
@@ 1.3. Structure of the paper
In Section˜2, we define the $D$-intersection map $\sigma_{\mathfrak{m}}$ in a local setting, namely for affine curves over discretely valued fields. We introduce $D$-transversal models, integral and rational reduction types $\Sigma_{\mathfrak{m}}$ and show that the composition $\sigma_{\mathfrak{m}}\circ\operatorname{AJ}_{P_{0}}$ restricted to points of a given reduction type lands in the subset $\mathfrak{S}_{\mathfrak{m}}(P_{0},\Sigma_{\mathfrak{m}})$. We also compare the $D$-intersection map with the local Kummer map. In Section˜3, we define the global $D$-intersection map $\sigma$ for an affine curve over a number field, and we determine the rank of $\ker(\sigma)$. We define $S$-integral reduction types $\Sigma$ and the global Selmer set $\operatorname{Sel}(P_{0},\Sigma)$. We then show that the Abel–Jacobi images of $S$-integral points of reduction type $\Sigma$ are contained in $\operatorname{Sel}(P_{0},\Sigma)$ and bound the rank of $\operatorname{Sel}(P_{0},\Sigma)$. In Section˜4, we summarise the theory of $p$-adic abelian integrals of log differentials. In Section˜5 we prove ˜A and describe an $S$-integral version of Siksek’s Restriction of Scalars Chabauty [Sik13]. We then proceed to prove ˜C in Section˜6, and end this paper in Section˜7 with three applications of our methods to even degree hyperelliptic curves: recovering the Linear Quadratic Chabauty method of [GM25], an improved bound on the number of integral points of such curves, and the complete determination of the $\mathbb{Z}[\zeta_{3}]$-points of a curve of genus 2 and Mordell–Weil rank 2.
@@ Acknowledgements
We thank Jennifer Balakrishnan, Alex Betts, Stevan Gajović, Sachi Hashimoto, Enis Kaya, Bjorn Poonen, Pim Spelier, and Elie Studnia for discussions on this paper. We thank Steffen Müller for helpful dicussions and comments on a previous version of this paper. The first author is supported by a Walter Benjamin Scholarship from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), project number LE 5634/1, and also acknowledges support from the DFG through TRR 326 Geometry and Arithmetic of Uniformized Structures, project number 444845124. The second author is supported by a Minerva Fellowship of the Minerva Stiftung Gesellschaft für die Forschung mbH and also acknowledges support through a guest postdoc fellowship at the Max Planck Institute for Mathematics in Bonn and through an NWO Grant, project number VI.Vidi.192.106.
@@ 2.2. Intersection numbers
For the remainder of this section, let $(R,\mathfrak{m},k)$ be a discrete valuation ring with field of fractions $K$, let $X$ be a smooth projective curve over $K$, let $D\neq\emptyset$ be a finite set of closed points of $X$ and set $Y=X\smallsetminus D$. In order to construct the $D$-intersection map $\sigma_{\mathfrak{m}}\colon J_{Y}(K)\to V_{D,\mathfrak{m}}$, we choose a regular model $\mathcal{X}$ of $X$ over $R$, i.e. an integral, regular, projective and flat $R$-scheme with an isomorphism $\mathcal{X}_{K}\cong X$ [Liu06, Definition 10.1.1]. We will show later that the map $\sigma_{\mathfrak{m}}$ does not depend on the chosen model. Recall from [Lan88, §III.2] the definition of intersection numbers on the arithmetic surface $\mathcal{X}$. If $E,F$ are effective divisors on $\mathcal{X}$ without common component and $x$ is a closed point of $\mathcal{X}_{\mathfrak{m}}$, the intersection number of $E$ and $F$ at $x$ is defined as
@@
$i_{\mathfrak{m}}(E,F)\coloneqq E\cdot F\coloneqq\deg_{k}(E.F)=\sum_{x}i_{x}(E,F)[k(x):k],$ (2.2)
where the sum is over all closed points of $\mathcal{X}_{\mathfrak{m}}$. By replacing a divisor with a linearly equivalent one, one can also define the self-intersection number $i_{\mathfrak{m}}(E,E)$ of a vertical prime divisor $E$. Hence, the symbol $i_{\mathfrak{m}}$ extends to all pairs of divisors on $\mathcal{X}$ with no horizontal common component [Lan88, §III.3].
@@ 2.6. $D$-transversal models
As $R$ is excellent, by [Liu06, Theorem 9.2.26] successive blow-ups of $\mathcal{X}$ in closed points [^1] As we only need $\operatorname{\mathcal{D}}$ to be regular at all $x\in\mathcal{X}_{\mathfrak{m}}^{\mathrm{sm}}(k)$, blow-ups in those points $x\in(\mathcal{X}_{\mathfrak{m}}^{\mathrm{sm}}\cap\operatorname{\mathcal{D}})(k)$ with $i_{x}(\mathcal{X}_{\mathfrak{m}},\operatorname{\mathcal{D}})\geq 2$ should suffice, but we did not check whether the analogue of [Liu06, Lemma 9.2.32] is true. yield a regular model $\mathcal{X}^{\prime}$ of $X$ in which the closure $\operatorname{\mathcal{D}}^{\prime}$ of $D$ is normal. Now we blow up again in every $x\in({\mathcal{X}^{\prime}_{\mathfrak{m}}}^{\!\mathrm{sm}}\cap\operatorname{\mathcal{D}}^{\prime})(k)$ such that $i_{x}(\mathcal{X}^{\prime}_{\mathfrak{m}},\operatorname{\mathcal{D}}^{\prime})\geq 2$ to obtain the model $\mathcal{X}^{\prime\prime}$. We claim that $\mathcal{X}^{\prime\prime}$ is a $D$-transversal regular model of $X$. Indeed, away from those $x$ the model $\mathcal{X}^{\prime}$ is already $D$-transversal. For $x\in({\mathcal{X}^{\prime}_{\mathfrak{m}}}^{\!\mathrm{sm}}\cap\operatorname{\mathcal{D}}^{\prime})(k)$, let $Q$ be the unique cusp such that $x$ lies in the closure $\mathcal{Q}^{\prime}$ of $Q$. If $2\leq i_{x}(\mathcal{X}^{\prime}_{\mathfrak{m}},\operatorname{\mathcal{D}}^{\prime})=i_{x}(\mathcal{X}^{\prime}_{\mathfrak{m}},\mathcal{Q}^{\prime})$, let $E$ denote the exceptional divisor on $\mathcal{X}^{\prime\prime}$ of the blow-up in $x$. Then $\mathcal{Q}^{\prime\prime}$ intersects $\mathcal{X}^{\prime\prime}_{\mathfrak{m}}$ in the intersection point of $E$ with the strict transform of ${\mathcal{X}^{\prime}_{\mathfrak{m}}}^{\!\mathrm{sm}}$, which is not a point in ${\mathcal{X}^{\prime\prime}_{\mathfrak{m}}}^{\mathrm{sm}}$, so there is no $D$-transversality left to check. ∎
@@ 3. Global theory: Selmer subsets of the generalised Jacobian
Let $K$ be a number field, let $X$ be a smooth projective curve over $K$, let $D\neq\emptyset$ be a finite set of closed points of $X$ and set $Y=X\smallsetminus D$. The ring of integers of $K$ is denoted by $\mathcal{O}_{K}$. We start by defining a $\mathbb{Q}$-vector space $V_{D}$ and a homomorphism $\sigma\colon J_{Y}(K)\to V_{D}$. In Section˜3.2 we define $S$-integral reduction types with respect to a regular model over $\mathcal{O}_{K}$ and use the map $\sigma$ to constrain the image of points of a fixed reduction type under an Abel–Jacobi map $Y\to J_{Y}$.
@@ 3.3. Selmer sets
Denote by $C(\Sigma)$ the set of primes $\mathfrak{q}\in S$ for which $\Sigma_{\mathfrak{q}}$ is cuspidal, i.e. an element of $\mathcal{C}_{\mathfrak{q}}^{1}=(\mathcal{X}_{\mathfrak{q}}^{\mathrm{sm}}\cap\operatorname{\mathcal{D}})(\operatorname{\mathbb{F}}_{\mathfrak{q}})$. The cuspidal part $\Sigma^{{\mathrm{csp}}}$ of $\Sigma$ is its image under the projection to $\prod_{\mathfrak{q}\in C(\Sigma)}\mathcal{C}_{\mathfrak{q}}^{1}$.
@@
Assume that $D$ does not consist of a single $K$-rational point. Then $\mathfrak{S}(P_{0},\Sigma)$ is a translate of a subgroup $U(\Sigma^{{\mathrm{csp}}})$ of $V_{D}$ which is free of rank $C(\Sigma)$.
@@
We compare the Selmer set $\operatorname{Sel}(P_{0},\Sigma)\subseteq J_{Y}(K)$ with the cohomological Selmer set $\operatorname{Sel}_{\Sigma,U_{1}}\subseteq\mathrm{H}^{1}(G_{K},V_{p}J_{Y})$ of [Bet23, §6.2], where $U_{1}=V_{p}J_{Y}$ is the abelianisation of the $\mathbb{Q}_{p}$-prounipotent fundamental group of $Y$. Assume that all primes $\mathfrak{q}|p$ are primes of good reduction for $(\mathcal{X},\operatorname{\mathcal{D}})$, that $P_{0}$ is integral at those primes, and that they are not contained in $S$. The set $\operatorname{Sel}_{\Sigma,U_{1}}$ is defined as the set of those global cohomology classes whose restriction at primes $\mathfrak{q}|p$ is crystalline, and whose restriction lies in $\mathfrak{S}^{\mathrm{coh}}_{\mathfrak{q}}(P_{0},\Sigma_{\mathfrak{q}})\subseteq\mathrm{H}^{1}(G_{K_{\mathfrak{q}}},V_{p}J_{Y})$ for all primes $\mathfrak{q}\nmid p$; see §2.8 for the definition of $\mathfrak{S}_{\mathfrak{q}}^{\mathrm{coh}}(P_{0},\Sigma_{\mathfrak{q}})$. ([Bet23, §6] also assumes that the base point $P_{0}$ is $S$-integral and that $K=\mathbb{Q}$ but the definition makes sense more generally.) Our assumptions ensure that there is only one reduction type $\Sigma_{\mathfrak{q}}$ at primes $\mathfrak{q}\mid p$ and $\mathfrak{S}_{\mathfrak{q}}(P_{0},\Sigma_{\mathfrak{q}})=\{0\}$. Let $\kappa\colon J_{Y}(K)\to\mathrm{H}^{1}(G_{K},V_{p}J_{Y})$ denote the global Kummer map, which is defined in the same way as in (2.16) and compatible with the local Kummer maps. Now ˜2.33 and ˜2.36 show